# Red-Teaming for Generative AI: Silver Bullet or Security Theater?

### v2

@@ 1 Introduction
In recent years, generative AI technologies, including large language models (LLMs) [146, 4] image and video generation models [115, 121, 21], and audio generation models [46, 5] have captured public imagination. While many view the proliferation and accessibility of these tools favorably, envisioning boons to productivity, creativity, and economic growth, concerns have emerged that the rapid adoption of these powerful models might unleash new categories of societal harms. These concerns have gained credibility owing to several well-publicized problematic incidents where such AI output text expressing discriminatory sentiment towards marginalized groups [93, 56, 101, 60, 62], created images reflecting harmful stereotypes [86, 151], and enabled the generation of deepfake audio in a fashion that has been likened to digital blackface [49]. These issues are compounded by the lack of transparency and accountability surrounding the development and evaluation of these models [4, 161, 16].
@@
On one hand, red-teaming appears to call for the right stuff: find the flaws, find the vulnerabilities, and (help to) eliminate them. In this spirit, one might find its inclusion in a landmark policy document to be welcome. On the other hand, for all of the virtue in its aims, red-teaming at this level of description is strikingly vague. As noted by The Frontier Model Forum (2023) [142], “there is currently a lack of clarity on how to define ‘AI red teaming’ and what approaches are considered part of the expanded role it plays in the AI development life cycle.” For example, the definition offered by the presidential executive order leaves the following key questions unanswered: What types of undesirable behaviors, limitations, and risks can or should be effectively caught and mitigated through red-teaming exercises? How should the activity be structured to maximize the likelihood of finding such flaws and vulnerabilities? For instance, aside from AI developers, who else should be at the table, and what resources should be available to them? How should the risks identified through red-teaming be documented, reported, and managed? Is red-teaming on its own sufficient for assessing and managing the safety, security, and trustworthiness of AI? If not, what other practices should be part of the broader evaluation toolbox, and how does red-teaming complement those approaches? In short, is red-teaming the stuff of policy—the sort of concrete practice around which we can structure regulatory requirements?; or is it the stuff of vibes—a vague practice better suited to rallying than to rule-making?
Methodology. Using publicly available resources, we gathered information about recent real-world cases of AI red-teaming exercises (Section 3). We emphasize that many of these cases stem from private sector companies who may utilize other evaluation practices and techniques not shared with the general public. As such, our corresponding analyses and conclusions rest on disclosed details. To complement these case studies primarily conducted by industry, we additionally performed an extensive survey of existing research literature on red-teaming and adjacent testing and evaluation methods (e.g., penetration testing, jailbreaking, and beyond) for generative AI (Section 4). We organized the thematic analysis of our case studies and literature survey around the following key questions:
Definition and scope: What is the working definition of red-teaming? What is the criterion for successful red-teaming?
@@
To extend and validate our analysis, we further analyzed public comments submitted to a Request For Information (RFI) issued by the National Institute of Standards and Technology (NIST) arm of the Department of Commerce. This RFI sought opinions on points relevant to red-teaming as outlined in the Executive Order. Finally, we conducted a workshop that convened leading experts in research, industry, and policy to discuss similar ideas and issues.111See appendix for takeaways from RFI analysis and workshop.
@@
Contributions. Our findings reveal a lack of consensus around the scope, structure, and assessment criteria for AI red-teaming. Prior methods and practices of AI red-teaming diverge along several critical axes, including the choice of threat model (if one is specified), the artifact under evaluation, the setting in which the activity is conducted (including actors, resources, and methodologies, and test-beds), and the resulting decisions the activity instigates (e.g., reporting, disclosure, and mitigation). In light of our findings, we argue that while red-teaming may be a valuable big-tent idea, and even a useful framing for a broad set of evaluation activities for generative AI models, the bludgeoning use of AI red-teaming (as defined in public literature) as a catch-all response to quiet all regulatory concerns about model safety verges on security theater [79]. Our work, featuring NIST RFI comment analysis and takeaways from an expert convening which both support our case studies and literature review, shows that the current framing of red-teaming in the public discourse serves more to assuage regulators and other concerned parties than to offer a concrete solution. To move toward a more robust toolbox of evaluations for generative AI, we synthesize our recommendations into a question bank meant to guide and scaffold future AI red-teaming practices (see Table 1) and propose future research, including improving the question bank through co-design and evaluation.
@@ 2 Related Contemporary Work
A brief history of red-teaming. Zenko [175] and Abbass [2] describe how the key concepts of red-teaming originated hundreds of years ago in warfare and religious contexts. They note the term “red team” was formally applied by the US military as early as the 1960s when modeling the Soviet Union’s behavior (in contrast to the “blue team” representing the US). According to Abbass et al. [1] and Abbass [2], red-teaming in computer security involves modeling an adversary and “map[ping] out the space of vulnerabilities from a threat lens” in contrast to penetration testing (in which enlisted cybersecurity experts actively attempt to find vulnerabilities in a computer system). Wood & Duggan [162] further describe how red-teaming “is not an audit” and that interpreting it as such risks reducing the amount of information shared about possible vulnerabilities. Using a hypothetical pandemic example, Bishop et al. [17] argue that effectively red-teaming a system requires context, knowledge, and assumptions about system usage.
Evaluation beyond red-teaming. Chang & Custis [30] note that red-teaming is only one of many approaches to increase transparency of an AI system and that factsheets, audits, and model cards are other ways to do so. Similarly, Horvitz [63] warns of more advanced deepfakes in the near future while emphasizing that remedies such as increased media literacy and output watermarking (that flags relevant media as AI-generated) should be employed alongside red-teaming; Kenthapadi et al. [69] echo these concerns and similar solutions in their tutorial. Shevlane et al. [135] also argue that both internal and external model evaluations, as well as robust security responses, should complement effective red-teaming to counter GenAI risks.
@@
3 Case Studies: Prior AI Red-teaming
@@
Methodology. We sourced case studies for this review by searching for reports and news stories about recent red-teaming exercises. As such, our selection is not meant to reflect the full range of red-teaming activities conducted in practice, as limited disclosures from industry teams would make such a reflection impossible. This said, the evaluations we cover here were mostly conducted by private companies, and they encompass a broad range of methods, goals, and areas of focus. In total, we surface and analyze six red-teaming exercises based on retrieved public reports. Table 2 contains more information and related sources pertaining to each of these exercises.
@@ 3.1 Findings
Variation in goals, processes, and threat models. Reflecting the lack of consensus on a definition of red-teaming in the literature, red-teaming activities frequently varied in form and in goals. Some organizations chose to conduct a single round of red-teaming [104, 8, 29, 53], while others saw red-teaming as an iterative process in which results from initial rounds of testing were used to prioritize risk areas for further investigation [142, 95, 4]. The goals of red-teaming activities also ranged from specific objectives (e.g., red-teaming to investigate risks to national security [8]) to more broad targets (e.g., uncovering “harmful” model behavior [142]); threat models related to the latter were more common. Model developers often use broader threat models for evaluation in hopes that this will yield greater variation in red-teaming efforts (especially because it is often impossible to fully understand the model’s entire risk surface or list all possible failure modes). Unfortunately, probing these nonspecific threat models does not always produce this desired variation, more so when evaluators are given limited time and resources to produce harmful model outputs. For example, some time-boxed evaluators repeatedly probed models in the same risk area because it was easy to produce harmful outputs, as opposed to further exploring the model’s full risk surface [53].
Interconnected members, resources, and outcomes. The evaluators employed in red-teaming activities for each case study varied considerably. Teams ranged from groups of subject matter experts to random samplings of community stakeholders. We found that there were generally three types of team compositions:
@@
The resources available to evaluation teams seemed to vary based on team composition. For crowdsourced teams, red-teaming efforts were time-boxed either by participant or by task, and access to models was available only through APIs [28, 53]. For teams with subject matter expertise, red-teaming efforts were more open-ended with fewer restrictions on time or compute [95, 4, 8]. While API access to models is still most common for these teams, sometimes experts are given access to versions of models without safety guardrails. When language models are used to red-team themselves, the main resource bottlenecks are the number of prompts used to produce red-teaming behavior and the compute resources needed for model retraining or fine tuning. Full access to model parameters is thus usually a requirement when performing this type of red-teaming [104]. As such, team composition and available resources also shape red-teaming outcomes. For instance, crowdsourced teams typically focused on risk areas where successful attacks were easy to produce due to time constraints, so risk areas that are more complex to attack may remain completely untested [53, 28]. In contrast, subject matter experts and members of academic communities and AI firms prioritized different risks and explored them in more detail due to differences in team member selection and resources [4]. When using language models for red-teaming, offensiveness classifiers are often trained on pre-existing datasets such as the Bot-Adversarial Dialogue (BAD) dataset [167], which in turn only cover certain types of offensive model replies. Evidently, team selection and resources can introduce bias into the types of risks investigated and ultimate exercise findings.
@@
No standards for disclosing red-teaming details. We found nontrivial variation in the publicly-shared outputs of red-teaming efforts, largely because there are no existing standardized procedures or requirements for reporting results of red-teaming. In only half of the cases explored, specific examples of risky or harmful model behavior uncovered by red-teaming efforts were publicly shared. In one case, a full dataset composed of 38,961 red team attacks was publicly released to aid in testing of other models [53]. In the other two cases, examples of harmful behavior were publicly available, but the full scope of all red-teaming attacks was not released [4, 104]. For red-teaming efforts on publicly available models or those focused on national security, specifics of harmful behavior were not shared publicly because findings were deemed “too sensitive.” One case study resulted in Anthropic piloting a responsible disclosure process to share vulnerabilities identified during red-teaming with appropriate community stakeholders, but this process is still under development (thus we assume that these disclosures have not yet been made) [8]. There are also variations in reporting on resource consumption. We found costs of red-teaming efforts were usually disclosed for evaluation teams composed of crowdsourced evaluators (for example, the hourly rate paid to crowdworkers [53]). These details were not disclosed for teams composed of subject matter experts and language models, though they seem to have been given greater time and compute resources. Two case studies specifically mention ongoing red-teaming for 6-7 months before model release [4, 9]. In contrast, for crowdsourced teams, evaluators spent about 30-50 minutes per task, with evaluators sourced from live events being limited to only completing a single task [53, 28].
@@
Diverse mitigations and supporting evaluations. While every case analyzed here identified problematic or risky model behavior, none of them resulted in a decision not to release the model. Instead, a number of risk mitigation strategies were proposed and/or employed to minimize harmful model behavior identified during red-teaming. These approaches ranged from concrete, such as jointly training language models and red-teaming models via strategies for training GANs, to purely conceptual, like unlikelihood training to reduce harmful outputs [104]. However, the specifics of risk mitigation strategies were often not provided when the target model was publicly available, and there were no standards for reporting improvements stemming from these efforts. As a result, it was often difficult to determine if risks identified during red-teaming were sufficiently addressed. Similarly, every case we analyzed involved models that had been previously evaluated using other techniques beyond red-teaming, but there were no guidelines or standards established for these evaluation methods. Commonly, models were evaluated using the Perspective API to measure toxicity; human feedback on helpfulness, harmfulness, and honesty; and QA benchmarks for accurate and truthful outputs [111, 9, 10]. Other evaluations included internal quantitative assessments to determine if model outputs violated specific content policies (e.g., hate speech, self-harm advice, illicit advice) [4]. Additionally, some initial efforts described as “red-teaming” by evaluators appeared more focused on understanding base model capabilities through open-ended experimentation than on specifically stress testing the model for vulnerabilities [95].
@@ 3.2 Discussion
Using notes collected from reports on each of the six case studies and the main points described above, two of the authors synthesized high-level takeaways from case study findings which are summarized below. Each author worked alone.
@@
Red-teaming is ill-structured. Evaluation teams seem aware that the entire risk surface of a model will not be explored by red-teaming activities. As such, they either prioritize risk areas for investigation or provide evaluators with broad directions in hopes that diversity within the group of evaluators will lead to the exploration of many different risks. However, in line with findings from prior empirical work [38, 43], there is a tradeoff between providing evaluators with specific instructions and exploring a variety of risk areas through red-teaming. On one hand, vague instructions can be helpful to avoid biasing evaluators towards finding specific issues based on initial prioritization. On the other, the lack of instructions could reduce the utility of the exercise for uncovering risks relevant to real-world contexts. We argue that this limited scope of red-teaming efforts is concerning. Namely, recent executive orders and evaluation frameworks establishing red-teaming as a best-practice suggest that the broader perception of red-teaming may not align with current working definitions of red-teaming, (i.e., red-teaming activities are much more qualitative, subjective and exploratory than community stakeholders may realize). In every case study, however, red-teaming was able to reveal harmful model behavior that other more systematic methods seemed to miss, highlighting the importance of both conducting red-teaming (alongside other evaluations) and developing systematic processes for red-teaming in a more comprehensive manner. These processes could include, for example, the development of guidelines on whether red-teaming is most effective when conducted internally or externally and when it should be conducted (i.e., before and/or after public release of the model and whether red-teaming activities should be ongoing while the model is publicly available).
@@
Evaluation team composition introduces biases. The goal of team member selection seems to be ensuring variety in the risk areas explored during red-teaming. One option used to get this variety is to handpick a team of experts with different backgrounds; another option is to use a random sampling of the population through crowdsourcing. Both of these options have drawbacks: there may be bias in the processes used to select experts [66, 35], and crowdworkers have very limited resources in terms of time, compute, and relevant expertise [147]. It is difficult to say what the ideal balance between expert and non-technical stakeholders would be, but prior crowdsourcing research suggests a hybrid approach could help address some of the pitfalls associated with each type of team composition [70, 150, 35]. One type of team composition we did not see explored in any case study is a crowdsourced team with more open-ended instructions and greater resources. This could allow more variety in the risk areas explored because evaluators would not feel incentivized to focus on risk areas where harmful model outputs are easy or quick to produce, but it would also require partnering with subject matter experts to fully evaluate risky model behavior. Team composition can also shape the outputs of red-teaming. One of the issues with red-teaming via internal teams is that more extreme measures such as blocking the release of a model will probably never be recommended due to conflicting interests. On the other hand, external teams that may be more likely to recommend such measures often do not have the power to actually employ these mitigations. Again, a hybrid approach could resolve some of these issues, but it would need to be paired with accountability mechanisms to disclose recommendations and mitigations.
@@
Hesitancy to release results reduces utility. As suggested by prior work studying responsible AI industry practices [88, 114], the reluctance to share all results from red-teaming activities may stem from risks associated with public models (evaluators do not want to provide inspiration for potential attackers). Additionally, releasing all of the data associated with red-teaming could be overwhelming for community stakeholders. This said, because red-teaming does not seem to be planned as a comprehensive measure of risky model behavior, disclosing some specifics of red-teaming efforts and resulting mitigation strategies is necessary so stakeholders can understand harms investigated and in turn determine if they are relevant to their use cases. For example, significant risk areas that evaluation teams knowingly have not probed should be highlighted or identified in reports. Moreover, none of the case studies provided complete monetary costs of red-teaming efforts. This information seems relatively low-risk to release (compared to specific examples of harmful model behavior, for instance) and could be useful for developing methods to conduct more comprehensive red-teaming. For example, the costs of assembling teams with differing compositions of experts, non-technical stakeholders, and automation could be used to determine where resources can be used most effectively, especially considering that a hybrid team composition could lead to the best coverage of the risk surface. Additionally, the costs of evaluating and mitigating various types of risks could be factored into a cost-benefit analysis when prioritizing risks based on real-world impact. The lack of reported cost figures may also make it harder for third-party or external organizations to conduct red-teaming: if these unreported costs are quite large, it could be difficult or impossible for anyone aside from companies themselves to do this type of analysis [39]. This information would be invaluable when developing guidelines for red-teaming to advise decisions such as whether internal or external red-teaming is most effective.

### v3

@@ 1 Introduction
In recent years, generative AI technologies, including large language models (LLMs) [154, 4] image and video generation models [121, 129, 22], and audio generation models [49, 5] have captured public imagination. While many view the proliferation and accessibility of these tools favorably, envisioning boons to productivity, creativity, and economic growth, concerns have emerged that the rapid adoption of these powerful models might unleash new categories of societal harms. These concerns have gained credibility owing to several well-publicized problematic incidents where such AI output text expressing discriminatory sentiment towards marginalized groups [97, 60, 106, 64, 66], created images reflecting harmful stereotypes [90, 159], and enabled the generation of deepfake audio in a fashion that has been likened to digital blackface [52]. These issues are compounded by the lack of transparency and accountability surrounding the creation of these models [17, 4, 169].
@@
On one hand, red-teaming appears to call for the right stuff: find the flaws, find the vulnerabilities, and (help to) eliminate them. In this spirit, one might find its inclusion in a landmark policy document to be welcome. On the other, for all of the virtue in its aims, red-teaming at this level of description is strikingly vague. As noted by The Frontier Model Forum (2023) [150], “there is currently a lack of clarity on how to define ‘AI red teaming’ and what approaches are considered part of the expanded role it plays in the AI development life cycle.” For example, the definition offered by the presidential executive order leaves the following key questions unanswered: What types of undesirable behaviors, limitations, and risks can or should be effectively caught and mitigated through red-teaming exercises? How should the activity be structured to maximize the likelihood of finding such flaws and vulnerabilities? For instance, aside from AI developers, who else should be at the table, and what resources should be available to them? How should the risks identified through red-teaming be documented, reported, and managed? Is red-teaming on its own sufficient for assessing and managing the safety, security, and trustworthiness of AI? If not, what other practices should be part of the broader evaluation toolbox, and how does red-teaming complement those approaches? In short, is red-teaming the stuff of policy—the sort of concrete practice around which we can structure regulatory requirements?; or is it the stuff of vibes—a vague practice better suited to rallying than to rule-making?
Methodology. Using publicly available resources, we gathered information about recent real-world cases of AI red-teaming exercises (Section 3). We emphasize that many of these cases stem from private sector companies who may use other evaluation techniques not shared with the general public. As such, our corresponding analyses and conclusions rest on disclosed details. To complement these case studies primarily conducted by industry, we additionally performed an extensive survey of existing research literature on red-teaming and adjacent testing and evaluation methods (e.g., penetration testing, jailbreaking, and beyond) for generative AI (Section 4). We organized the thematic analysis of our case studies and literature survey around the following key questions:
Definition and scope: What is the working definition of red-teaming? What is the success criterion?
@@
To further extend and validate our analysis, we analyzed public comments submitted to a Request For Information (RFI) issued by the National Institute of Standards and Technology (NIST) arm of the Department of Commerce. This RFI sought opinions on points relevant to red-teaming as outlined in the Executive Order.111See appendix for RFI comment analysis.
@@
Contributions. Our findings reveal a lack of consensus around the scope, structure, and assessment criteria for AI red-teaming. Prior methods and practices of AI red-teaming diverge along several critical axes, including the choice of threat model (if one is specified), the artifact under evaluation, the setting in which the activity is conducted (including actors, resources, methodologies, and test-beds), and the resulting decisions the activity instigates (e.g., reporting, disclosure, and mitigation). In light of our findings, we argue that while red-teaming may be a valuable big-tent idea, and even a useful framing for a broad set of evaluation activities for generative AI models, the bludgeoning use of AI red-teaming (as defined in public literature) as a catch-all response to quiet all regulatory concerns about model safety verges on security theater [83]. Our work, including our NIST RFI comment analysis, shows that the current framing of red-teaming in the public discourse serves more to assuage regulators and other concerned parties than to offer a concrete solution. To move toward a more robust toolbox of evaluations for generative AI, we synthesize our recommendations into a question bank meant to guide and scaffold future AI red-teaming practices (see Table 1) and propose future research, including improving the question bank through co-design and evaluation.
@@ 2 Related Contemporary Work
A brief history of red-teaming. Zenko [184] and Abbass [2] describe how the key concepts of red-teaming originated hundreds of years ago in warfare and religious contexts. They note the term “red team” was formally applied by the US military as early as the 1960s when modeling the Soviet Union’s behavior (in contrast to the “blue team” representing the US). In computer security, red-teaming involves modeling an adversary and “map[ping] out the space of vulnerabilities from a threat lens” in contrast to penetration testing (in which enlisted cybersecurity experts actively attempt to find vulnerabilities in a computer system) [1, 2]. Wood & Duggan [170] further describe how red-teaming “is not an audit” and that interpreting it as such risks reducing the amount of information shared about possible vulnerabilities. Using a hypothetical pandemic example, Bishop et al. [18] argue that effectively red-teaming a system requires context, knowledge, and assumptions about system usage.
Evaluation beyond red-teaming. Chang & Custis [31] note that red-teaming is only one of many approaches to increase transparency of an AI system and that factsheets, audits, and model cards are other ways to do so. Similarly, Horvitz [67] warns of more advanced deepfakes in the near future while emphasizing that remedies such as increased media literacy and output watermarking (flagging relevant media as AI-generated) should be employed alongside red-teaming; Kenthapadi et al. [73] echo these concerns and similar solutions in their tutorial. Shevlane et al. [143] also argue that both internal and external model evaluations, as well as robust security responses, should complement effective red-teaming to counter GenAI risks.
@@
3 Case Studies: AI Red-teaming in practice
@@
Methodology. We sourced case studies by searching for reports and news stories about recent red-teaming exercises. As such, our selection is not meant to reflect the full range of red-teaming activities conducted in practice, as limited disclosures from industry teams would make such a reflection impossible. This said, the evaluations we cover here were mostly conducted by private companies, and they encompass a broad range of methods, goals, and areas of focus. In total, we surface and analyze six red-teaming exercises based on retrieved public reports. See Table 2 for more information.
@@ 3.1 Findings
Variation in goals, processes, and threat models. Reflecting the lack of consensus on a definition of red-teaming in the literature, red-teaming activities frequently varied in form and in goals. Some organizations chose to conduct a single round of red-teaming [57, 109, 8, 30], while others saw red-teaming as an iterative process in which results from initial rounds of testing were used to prioritize risk areas for further investigation [4, 150, 99]. The goals of red-teaming activities also ranged from specific objectives (e.g., red-teaming to investigate risks to national security [8]) to more broad targets (e.g., uncovering “harmful” model behavior [150]); threat models related to the latter were more common. Model developers use such threat models for evaluation in hopes that this will yield greater variation in red-teaming efforts, especially because it is impossible to understand the model’s entire risk surface. Unfortunately, probing these nonspecific threat models does not always produce this desired variation, more so when evaluators are given limited time and resources to produce harmful outputs. For example, some time-boxed evaluators repeatedly probed models in the same risk area because it was easy to produce harmful outputs, as opposed to exploring other risks [57].
Interconnected members, resources, and outcomes. The evaluators employed in red-teaming activities for each case study varied considerably. We found that there were generally three types of team compositions:
@@
The resources available to evaluation teams varied based on team composition. For crowdsourced teams, red-teaming efforts were time-boxed either by participant or by task, and access to models was available only through APIs [57, 29]. For teams with subject matter expertise, red-teaming efforts were more open-ended with fewer restrictions on time or compute [4, 8, 99]. While API access to models is still most common for these teams, sometimes experts are given access to versions of models without safety guardrails. When language models are used to red-team themselves, the main resource bottlenecks are the number of prompts used to produce red-teaming behavior and the compute resources needed for model retraining or fine tuning. Full access to model parameters is thus usually a requirement when performing this type of red-teaming [109]. As such, team composition and available resources also shape red-teaming outcomes. For instance, crowdsourced teams typically focused on risk areas where successful attacks were easy to produce due to time constraints, so risk areas that are more complex to attack may remain completely untested [57, 29]. In contrast, subject matter experts and members of academic communities and AI firms prioritized different risks and explored them in more detail due to differences in team member selection and resources [4]. When using language models for red-teaming, offensiveness classifiers are trained on pre-existing datasets such as the Bot-Adversarial Dialogue (BAD) dataset [176], which in turn only cover certain types of offensive model replies. Evidently, team selection and resources can introduce bias into the types of risks investigated and ultimate exercise findings.
@@
No standards for disclosing red-teaming details. We found nontrivial variation in the publicly-shared outputs of red-teaming efforts, largely because there are no existing standardized reporting procedures or requirements. In only half of the cases explored, specific examples of risky or harmful model behavior uncovered by red-teaming efforts were publicly shared. In one case, a full dataset composed of 38,961 red team attacks was publicly released to aid in testing of other models [57]. In the other two cases, examples of harmful behavior were publicly available, but the full scope of all red-teaming attacks was not released [4, 109]. For red-teaming efforts on publicly available models or those focused on national security, specifics of harmful behavior were not shared publicly because findings were deemed too sensitive to share without responsible disclosure practices [8]. One case study resulted in Anthropic piloting a responsible disclosure process to share vulnerabilities identified during red-teaming with appropriate community stakeholders, but this process is still under development (thus we assume that these disclosures have not yet been made) [8]. There are also variations in reporting on resource consumption. We found costs of red-teaming efforts were usually disclosed for evaluation teams composed of crowdsourced evaluators (for example, the hourly rate paid to crowdworkers [57]). These details were not disclosed for teams composed of subject matter experts and language models, though they seem to have been given greater time and compute resources. Two case studies specifically mention ongoing red-teaming for 6-7 months before model release [4, 9]. In contrast, for crowdsourced teams, evaluators spent about 30-50 minutes per task, with evaluators sourced from live events being limited to only completing a single task [57, 29].
@@
Diverse mitigations and supporting evaluations. While every case analyzed here identified problematic or risky model behavior, none of them resulted in a decision not to release the model. Instead, a number of risk mitigation strategies were proposed and/or employed to minimize harmful model behavior identified during red-teaming. These approaches ranged from concrete, such as jointly training language models and red-teaming models via strategies for training GANs, to purely conceptual, like unlikelihood training to reduce harmful outputs [109]. However, the specifics of risk mitigation strategies were often not provided when the target model was publicly available, and there were no standards for reporting improvements stemming from these efforts. As a result, it was often difficult to determine if risks identified during red-teaming were sufficiently addressed. Similarly, every case we analyzed involved models that had been previously evaluated using other techniques beyond red-teaming, but there were no established guidelines or standards for these other methods. Commonly, models were evaluated using the Perspective API to measure toxicity; human feedback on helpfulness, harmfulness, and honesty; and QA benchmarks for accurate and truthful outputs [117, 11, 9]. Other evaluations included internal quantitative assessments to determine if model outputs violated specific content policies (e.g., hate speech, illicit advice) [4]. Additionally, some initial efforts described as “red-teaming” by evaluators were more focused on understanding base model capabilities through open-ended experimentation than on specifically stress testing the model [99].
@@ 3.2 Discussion
Red-teaming is ill-structured. Evaluation teams either prioritize risk areas for investigation or provide evaluators with broad directions in hopes that diversity within the group of evaluators will lead to the exploration of many different risks. However, in line with findings from prior empirical work [39, 45], there is a tradeoff between providing evaluators with specific instructions and leaving the activity open-ended. On one hand, vague instructions can be helpful to avoid biasing evaluators towards finding specific issues based on initial prioritization. On the other, a lack of instructions can reduce the utility of the exercise for uncovering risks relevant to real-world contexts. Red teams navigate this tradeoff as they seem aware that the entire risk surface of a model will not be explored by red-teaming activities, but this serves to make red-teaming as a whole ill-structured and difficult to define. Moreover, we argue that this lack of structure and scope is concerning as recent recommendations establishing red-teaming as a best practice suggest that the broader perception of red-teaming may not align with current working definitions of red-teaming, (i.e., red-teaming activities are much more qualitative, subjective and exploratory than community stakeholders may realize). In every case study, however, red-teaming was able to reveal harmful model behavior that other more systematic methods seemed to miss, highlighting the importance of both conducting red-teaming (alongside other evaluations) and developing systematic processes for red-teaming in a more comprehensive manner. These processes could include, for example, the development of guidelines on whether red-teaming is most effective when conducted internally or externally and when it should be conducted (i.e., before and/or after public release of the model and whether red-teaming activities should be ongoing while the model is publicly available).
@@
Evaluation team composition introduces biases. The goal of team member selection seems to be ensuring variety in the risk areas explored during red-teaming. One way to do so is by handpicking experts with different backgrounds, as noted by prior work on interdisciplinary collaboration within AI teams [102, 46]; another is by randomly sampling the population through crowdsourcing. Both have drawbacks: there may be bias in expert selection [70, 36], and crowdworkers have limited resources in terms of time, compute, and relevant expertise [155]. It is difficult to say what the ideal balance between expert and non-technical stakeholders would be, but prior crowdsourcing research suggests a hybrid approach could help address some of the pitfalls associated with each type of team composition [74, 158, 36]. One type of team composition we did not see explored in any case study is a crowdsourced team with more open-ended instructions and greater resources. This could allow more variety in the risk areas explored because evaluators would not feel incentivized to focus on risk areas where harmful model outputs are easy or quick to produce, but it would also require partnering with subject matter experts to fully evaluate risky model behavior. Team composition can also shape the outputs of red-teaming. One of the issues with red-teaming via internal teams is that more extreme measures such as blocking the release of a model may never be recommended due to conflicting interests. However, external teams that may be more likely to recommend such measures often do not have the power to actually employ these mitigations. A hybrid approach could resolve some of these issues, but it would need to be paired with accountability mechanisms to disclose recommendations and mitigations. Additionally, directly involving marginalized stakeholders, as they may suffer the most from unanticipated model outputs, is challenging yet not impossible (see, e.g., [171]).
@@
Hesitancy to release results reduces utility. As suggested by prior work studying responsible AI industry practices [92, 120], the reluctance to share all results from red-teaming activities may stem from risks associated with public models (evaluators do not want to provide inspiration for potential attackers). Additionally, releasing all of the data associated with red-teaming could be overwhelming for community stakeholders. This said, because red-teaming does not seem to be planned as a comprehensive measure of risky model behavior, disclosing some specifics of the activity is necessary so stakeholders can understand harms investigated and in turn determine if they are relevant to their use cases. For example, significant risk areas that evaluation teams knowingly have not probed should be highlighted or identified in reports. Moreover, none of the case studies provided complete monetary costs of red-teaming efforts. This information seems relatively low-risk to release (i.e., compared to specific examples of harmful model behavior) and could be useful for developing methods to conduct more comprehensive red-teaming. The costs of assembling teams with differing compositions of expertise and automation, for instance, could be used to determine where resources can be used most effectively. Similarly, the costs of evaluating and mitigating various types of risks could be factored into a cost-benefit analysis when prioritizing risks. The lack of reported cost figures may also make it harder for third-party or external organizations to conduct red-teaming: if these unreported costs are quite large, it could be difficult or impossible for anyone aside from companies themselves to do this type of analysis [40].